Profitability and Asset Management Ratios You are thinking of investing in Tikki's Torches, Inc. You have only the following information on the firm at year-end 2008: net income = $680,000, total debt = $13.8 million, debt ratio = 43%. What is Tikki's ROE for 2008?
Answer #1

debt ratio=debt/total assets

total assets=(13.8/0.43) million

=$32.0930233 million

total assets=debt+equity

equity=(32.0930233-13.8)=$18.2930233 million(Approx)

=$18293023.3

ROE=net income/equity

=680,000/18293023.3

=3.72%(Approx).